Atta Kay Physics:
Atta Kay's Special Mock-----5/8/2018
A particle P is let fall vertically from a height to the ground. Another particle Q is projected vertically down 3s later with a velocity of 3m/s at an acceleration of 12m/s^2 from the same height. If both particles reached the ground simultaneously, calculate the:
(a) Time of each particle to reach the ground
(b) Distance traveled by each particle to the ground. Take (g=10m/s^2).
GOOD LUCK
1
Expert's answer
2018-08-06T10:55:08-0400
A particle P:
h=(gt^2)/2=5t_1^2
A particle Q:
h=v_0 t+(at^2)/2=3t+6t^2
t_1=t+3
Hence,
5〖(t+3)〗^2=3t+6t^2
Solve this equation
t=28.5 sec
t+3=31.5 sec
h=4961.25 metres

(a) Time of each particle to reach the ground
A particle P: t=31.5 sec
A particle Q: t=28.5 sec
(b) Distance traveled by each particle to the ground.
h=4961.25 metres